OpenStudy (otonogold):

I am confused. The answer askes What is the graph of the absolute value equation? y= -2|x| +5 So would I put the -2 in the absolute value brackets and it will be positive pointing up or will the graph still be negative pointing down?

OpenStudy (anonymous):

The function will be y=-2*x+5 for (x>0) and y=-2*(-x) for (x<0) and y=-2*(0) for (x=0).

OpenStudy (otonogold):

so it'll be -2x+5 ? @Noliec

OpenStudy (anonymous):

That depends on your x, think of it graphically. :D

OpenStudy (anonymous):

I was being careless on the last one btw, when x=0, y=-2*(0)+5 of course.

OpenStudy (anonymous):

So you have that when x=0 --> y=5 when x is GREATER than 0, the function y behaves like y(x)=-2*x+5. when x is LESS than 0, the function behaves like y(x)=-2*(-x)+5.

OpenStudy (anonymous):

Think of it as plotting two graphs on each side of the y-axis, on the left side (x<0) so you have the linear equation I wrote above; on the right side (x>0) you have it behaving like y(x)=-2*x+5.

OpenStudy (otonogold):

So will it be pointing down since it is has a negative number or would it point up because it is greater than @noliec

OpenStudy (otonogold):

So will it be pointing down since it is has a negative number or would it point up because it is greater than @noliec

OpenStudy (phi):

y= -2|x| +5 So would I put the -2 in the absolute value brackets No, do not put the -2 inside the absolute value.... that changes the problem

OpenStudy (phi):

to find what this looks like, try a few small numbers for x try x= -2, -1, 0, 1, and +2 when x= -2 you get -2* |-2| + 5 = -2*2+ 5 = -4+5= 1 (-2,1) is a point on the graph when x = -1 you get y = 3 (-1,3) when x=0, you get y=5 (0,5) when x=+1 , y= 3 (1,3) now see which plot has those points